Use the contingency table to determine the following probabilities.

a. P( B or C)

b. P(B or D)

c. P(A or E)

d. P(A or B)

Event A

Event B

Event C

1111

1111

Event D

77

1010

Event E

1010

11

a.

P(B or C)=

​(Round to two decimal places as​ needed.)

b

P(B or D)=

​(Round to two decimal places as​ needed.)

c.

P(A or E)=

​(Round to two decimal places as​ needed.)

d.

P(A or B)=

​(Round to two decimal places as​ needed.)
Answer #1

Here we have :

Total data = 1111 + 1111 + 77 + 1010 + 1010 + 11 = 4330

Hence,

a) P(B or C ) = (1111 + 1010 + 11 + 1111) / 4330 = 0.75

b) P(B or D) = (1111 + 1010 + 11 + 77) / 4330 = 0.51

c) P(A or E) = (1111 + 77 +1010 +11 ) / 4330 = 0.51

d) P(A or B) = 4330/4330 = 1